# Exploring End-to-end Differentiable Neural Charged Particle Tracking – A Loss Landscape Perspective

### v1

@@ 4.2 Edge Scoring Architecture
Following the basic idea of track reconstruction using an edge classification scheme with edge scores (Heintz et al., 2020; DeZoort et al., 2021), we compose a graph neural network based on the interaction network (IN) architecture (Battaglia et al., 2016; 2018) as proposed in (DeZoort et al., 2021). However, we use a slightly modified architecture to predict node scores on the line graph representation $\mathcal{G}_{L}$, which directly correspond to the edge scores in $\mathcal{G}_{H}$ (see Figure 3). Following is the formulation of edge and node updates in a generic message passing formulation, as well as the final scoring function based on node representation on $\mathcal{G}_{L}$. In a first step, updated edge representations are calculated using a concatenated vector containing the edge attributes, describing the scattering behavior of the triplet-candidate, together with the node attributes $\bm{v}_{i}^{(0)}$ and $\bm{v}_{j}^{(0)}$ of adjacent nodes according to
@@
Here $\phi_{R,1}$ is a multilayer feed-forward network mapping the concatenated input to a fixed size representation with $\phi_{R,1}:\mathbb{R}^{2d_{\text{node}}+d_{\text{edge}}}\rightarrow\mathbb{R}^{d_{\text{hidden}}}$.
In a subsequent step, all node features of $\mathcal{G}_{L}$ are updated in an aggregation step using the current node representation and edge information aggregated from the direct neighborhood of the respective node $v_{i}$:
$\bm{v}_{i}^{\;(1)}=\phi_{O}\left(\bm{v}_{i}^{\;(0)},\max_{j\in\mathcal{N}_{i}}\left(\bm{e}_{ij}^{\;(1)}\right)\right)$ (10)
Given the dynamic number of incoming edges for each graph vertex, either due to the changing numbers of hit readouts in subsequent layers of $\mathcal{G}_{H}$ or the required generalization ability to different particle densities of the proton beam, we use a node-agnostic max-aggregation scheme, providing empirically better generalization abilities compared to non-node-agnostic aggregation schemes (Joshi et al., 2021).
To predict for each edge in $\mathcal{G}_{H}$ an edge score, we perform a final reasoning step, transforming the node embeddings of each node in $\mathcal{G}_{L}$ into a single scalar output using a feedforward network $\phi_{R,2}$:
@@
We generate the final edge scores $s_{ij}\in[0,1]$ using a sigmoid activation function as proposed in Heintz et al. (2020); DeZoort et al. (2021). We augment the multi-layer feed forward networks using batch normalization (Ioffe & Szegedy, 2015), to mitigate the risk of vanishing gradients and smoothen the optimization landscape (Santurkar et al., 2018), providing better training and inference performance.
@@ 4.3 Track Construction as a Differentiable Assignment Problem
Generating discrete track candidates from the continuous edge-score predictions, requires the application of a discrete algorithm that matches the tracks with the lowest total cost. Optimally, this would require a multidimensional assignment problem (MAP) to minimize the total cost of all assignments over the entire detector geometry. This exact definition is, however, impractical in real-world usage since solving a MAP is NP-hard (Pierskalla, 1968), requiring an approximate solution to solve. In the related literature, different optimization procedures, such as DBSCAN or union-find or partitioning algorithms, are used frequently (DeZoort et al., 2021; Ju et al., 2021; Lieret & DeZoort, 2023). In this work, however, we rely on a relaxation of the original MAP formulation to a layer-wise linear sum assignment problem (LSAP), finding a solution $y(\bm{\mathcal{C}})$ that minimizes
$\begin{split}y(\bm{\mathcal{C}}):\quad\min&\quad\sum_{(i,j)\in\mathcal{E}}\pi_{ij}c_{ij}\\ \text{s.t.}&\quad\sum_{i\in\mathcal{V}_{F}}\pi_{ij}=1,\quad j\in\mathcal{V}_{T}\\ &\quad\sum_{j\in\mathcal{V}_{T}}\pi_{ij}\leq 1,\quad i\in\mathcal{V}_{S}\\ &\quad\pi\in\{0,1\},\quad i,j\in\mathcal{V}_{T}\times\mathcal{V}_{S}\end{split}$ (12)
for every detector layer. Here, $c_{ij}\in\bm{\mathcal{C}}$ is the assignment cost for the edge combination $i,j$, where $i\in\mathcal{V}_{S}$ is a node from the set of source nodes and $j\in\mathcal{V}_{T}$ is a node from the set of target nodes respectively. Further, $\pi_{ij}$ is the respective assignment policy defining whether the edge should be assigned under the reconstruction policy. While the results in Section 5.1 demonstrate good performance for the choice of the LSA assignment, we want to note here that an increasing amount of detector noise and particle scattering can significantly influence the performance due to the required assignment without any notion of noise.
@@
For this general formulation of LSAP a wide range of work of predict-and-optimize schemes providing an end-to-end optimization ability to combinatorial optimization problems exist based on e.g., the interpolation of optimization mappings (Vlastelica et al., 2020; Sahoo et al., 2023), continuous relaxations (Amos & Kolter, 2017; Elmachtoub & Grigas, 2017; Wilder et al., 2019) or methods that bypass the calculation of gradients for the optimizer entirely using surrogate losses (Mulamba et al., 2021; Shah et al., 2022). For an in-depth review and comparison of existing approaches, the mindful reader is referred to Geng et al. (2023). Geng et al. (2023) demonstrated for all major and representative types of end-to-end training mechanisms for bipartite matching problems, similar regret bounds. We thus base our selection of a technique solely on simplicity and generalizability of the approach to other optimizers and use-cases. Vlastelica et al. (2020) defines a general blackbox differentiation scheme for combinatorial solvers of form $y(\bm{\mathcal{C}})=\arg\min_{y\in\mathcal{Y}}c(\bm{\mathcal{C}},y)$ by considering the linearization of the solver mapping at the point $y(\hat{\bm{\mathcal{C}}})$ according to
@@
Here, $y(\hat{\bm{\mathcal{C}}})$ and $y_{\lambda}(\bm{\mathcal{C}}^{\prime})$ are standard and perturbed-cost output of the combinatorial solver, $\lambda$ is a hyperparameter controlling the interpolation and $dL/dy$ is the gradient of the reconstruction loss w.r.t. the output of the combinatorial solver. Using a linearized view of the optimization mapping around the input allows remaining with exact solvers without the necessity of using any relaxation of the combinatorial problem.
@@
The usage of discrete assignment does not provide any confidence scores, thus task losses of correctly assigned edges are always exactly zero, even for marginal differences between cost elements, potentially having a noticeable impact on the generalization performance Rolínek et al. (2020b; a). To enforce larger margins between predicted costs, Rolínek et al. (2020b; a) introduced ground-truth-induced margins where a negative or positive penalty is added on the predicted costs based on the ground truth labels. Later, Sahoo et al. (2023) introduced noise-induced margins, removing the previous ground-truth dependence by adding random noise to the predicted costs. However, for the large number of edges in our use case of particle tracking, we found both methods to be highly unstable, even for small margin factors.
@@ 4.4 Network Optimization
($\mathcal{L}(\{y_{n}\}_{1:N},\{f(\mathcal{G}_{n})\}_{1:N})=\frac{1}{N}\sum_{n=1}^{N}\frac{1}{|\mathcal{E}_{n}|}\sum_{(i,j)\in\mathcal{E}_{n}}w_{ij}\cdot\left(y_{ij}\cdot(1-\hat{y}_{ij})+(1-y_{ij})\cdot\hat{y}_{ij}\right),$ (15)
@@ 5 Experimental Results and Analysis
For the studies reported in this work, we rely on MC simulation of detector readout data, which we generate using GATE (Jan et al., 2004; 2011) version 9.2 based on Geant4 (version 11.0.0) (Agostinelli et al., 2003; Allison et al., 2006; 2016). We generate different datasets for training (100,000 particles) and validation (5,000 particles), using a 230 MeV pencil beam as a beam source with water phantoms of various thicknesses between particle beam and detector. For comparability purposes, the test set, containing 10,000 particles in total, is taken from Kortus et al. (2022). A detailed listing of all data sources is provided in Appendix E. We generate for the training and validation simulations hit- and line graphs with 100 primaries per frame ($p^{+}/F$), according to the description in Section 4.1, which we consider the expected target density for the constructed detector. Further, we generate hit graphs for the test set with 50, 100 and 150 $p^{+}/F$ to assess the generalization ability to unseen densities.
@@
As baseline models, providing reference values for quantifying the track reconstruction performance of the proposed architectures, we select both a traditional iterative track follower algorithm (Pettersen et al., 2020) [^5] The source code for the track follower proposed by Pettersen et al. is provided as a software component in the Digital Tracking Calorimeter Toolkit (https://github.com/HelgeEgil/DigitalTrackingCalorimeterToolkit), and a reinforcement learning (RL) based tracking algorithm (Kortus et al., 2023). The first baseline finds suitable tracks by iteratively searching for track candidates minimizing the total scattering angle. Similarly, the RL-based algorithm aims to learn a reconstruction policy functioning as a learned heuristic to the algorithm by (Pettersen et al., 2020) by iteratively maximizing the expected likelihood of observing the scattering behavior determined by MCS during training.
@@ 5.1 Training and Inference Performance
Additional performance results are provided in Table 1. Here, the performance of the trained models is compared for different phantom and particle density configurations. In addition, the performance results of a traditional track follower procedure, which was developed for the particular use-case of particle tracking in the DTC prototype, are provided as a baseline. Similar to the results provided in Figure 4, all model configurations, trained on the task- and the prediction-loss, demonstrate near identical reconstruction purity and efficiency over all tested phantom and particle configurations. Further, the graph neural networks outperform the baseline algorithm in almost all performance metrics over all configurations, with only the configuration of 150 $p^{+}/F$ and 100 mm water phantom as a slight outlier in this tendency. Here, the track follower and RL-based tracker demonstrate a higher reconstruction efficiency, while the reconstruction purity is still higher for the GNN architectures.
Figure 4: True positive and false positive rate of assignments, as well as purity and efficiency of track reconstruction, evaluated for predict-and-track and predict-then-track on validation set as a function of training steps.
Table 1: Reconstruction performance, measured in terms of purity $p$ and efficiency $\epsilon$ for water phantoms of 100, 150 and 200 mm thickness and 100, 150 and 200 primaries per frame $p^{+}/F$. Results marked with * us a layer-wise reconstruction scheme with the initial transition generated using ground truth. Results for RL-based are taken from (Kortus et al., 2023)($\uparrow$: higher is better; $\downarrow$ lower is better)
@@
50 Track follower Pettersen et al. (2020) 87.1$\pm$0.0 78.1$\pm$0.0 89.4$\pm$0.0 81.1$\pm$0.0 90.8$\pm$0.0 82.1$\pm$0.0
RL-based Kortus et al. (2023)* 92.5$\pm$0.2 81.5$\pm$0.3 93.7$\pm$0.2 84.0$\pm$0.4 94.4$\pm$0.2 85.4$\pm$0.4
$\text{GNN}_{PTT}$ (BCE. edge score) 94.9$\pm$0.1 83.8$\pm$0.0 96.2$\pm$0.2 86.9$\pm$0.2 96.4$\pm$0.0 88.7$\pm$0.1
$\text{GNN}_{PAT}$ ($\lambda=25.0$) 94.9$\pm$0.1 83.8$\pm$0.1 96.2$\pm$0.1 87.0$\pm$0.1 96.5$\pm$0.1 88.9$\pm$0.1
$\text{GNN}_{PAT}$ ($\lambda=50.0$) 95.0$\pm$0.2 83.9$\pm$0.2 96.3$\pm$0.2 87.1$\pm$0.2 96.5$\pm$0.1 89.0$\pm$0.1
$\text{GNN}_{PAT}$ ($\lambda=75.0$) 95.0$\pm$0.2 83.9$\pm$0.2 96.3$\pm$0.1 87.1$\pm$0.1 96.5$\pm$0.0 89.0$\pm$0.0
100 Track follower Pettersen et al. (2020) 80.6$\pm$0.0 71.7$\pm$0.0 84.7$\pm$0.0 76.4$\pm$0.0 85.8$\pm$0.0 77.5$\pm$0.0
RL-based Kortus et al. (2023)* 85.6$\pm$0.3 75.2$\pm$0.5 88.8$\pm$0.5 79.0$\pm$0.5 89.5$\pm$0.4 80.8$\pm$0.5
$\text{GNN}_{PTT}$ (BCE. edge score) 87.4$\pm$0.3 75.2$\pm$0.2 91.9$\pm$0.3 82.4$\pm$0.3 91.7$\pm$0.2 83.8$\pm$0.1
$\text{GNN}_{PAT}$ ($\lambda=25.0$) 87.3$\pm$0.2 75.0$\pm$0.2 91.7$\pm$0.2 82.1$\pm$0.3 92.2$\pm$0.2 84.4$\pm$0.2
$\text{GNN}_{PAT}$ ($\lambda=50.0$) 87.4$\pm$0.3 75.1$\pm$0.2 92.0$\pm$0.1 82.4$\pm$0.2 92.5$\pm$0.1 84.6$\pm$0.1
$\text{GNN}_{PAT}$ ($\lambda=75.0$) 87.4$\pm$0.2 75.1$\pm$0.2 91.9$\pm$0.1 82.4$\pm$0.1 92.3$\pm$0.2 84.4$\pm$0.2
150 Track follower Pettersen et al. (2020) 75.6$\pm$0.0 67.2$\pm$0.0 80.1$\pm$0.1 72.2$\pm$0.0 82.5$\pm$0.0 74.6$\pm$0.0
RL-based Kortus et al. (2023)* 80.5$\pm$0.4 70.8$\pm$0.6 83.8$\pm$0.7 74.4$\pm$0.6 85.3$\pm$0.6 76.9$\pm$0.5
$\text{GNN}_{PTT}$ (BCE. edge score) 77.5$\pm$0.4 65.0$\pm$0.3 84.9$\pm$0.3 75.3$\pm$0.3 87.2$\pm$0.2 79.6$\pm$0.2
$\text{GNN}_{PAT}$ ($\lambda=25.0$) 76.7$\pm$0.2 64.3$\pm$0.2 84.8$\pm$0.3 75.1$\pm$0.3 87.6$\pm$0.3 80.1$\pm$0.3
$\text{GNN}_{PAT}$ ($\lambda=50.0$) 76.8$\pm$0.3 64.4$\pm$0.3 85.1$\pm$0.2 75.5$\pm$0.2 88.1$\pm$0.4 80.6$\pm$0.4
$\text{GNN}_{PAT}$ ($\lambda=75.0$) 76.6$\pm$0.1 64.3$\pm$0.1 85.0$\pm$0.2 75.4$\pm$0.2 87.8$\pm$0.2 80.4$\pm$0.2
@@ 5.2 Evaluation of Local and Global Loss Landscapes
To analyze and characterize the training and inference behavior of the end-to-end and two-step tracking approaches and gain an understanding of the effects of end-to-end optimization and its hyperparameters, we visualize (Li et al., 2018) and characterize the loss landscape structure, closely following the methodology and taxonomy by Yang et al. (2021) for characterizing the optimization loss landscapes. We specifically focus on the global connectivity of the loss landscape as it allows us to infer conclusions and comparisons between different model types, given the same parameterization along all configurations. We thus further augment the evaluation by additional analysis of functional similarities (Klabunde et al., 2023) of the optimized models. All used algorithms are detailed in Section 3.2. Additional implementation details are listed in Appendix D.

### v2

@@ 4.2 Edge Scoring Architecture
Following the basic idea of track reconstruction using an edge classification scheme (Heintz et al., 2020; DeZoort et al., 2021), we compose a graph neural network based on the interaction network (IN) architecture (Battaglia et al., 2016; 2018) as proposed in (DeZoort et al., 2021), predicting edge scores capturing the likelihood of every possible track segment candidate in $\mathcal{G}_{H}$. However, we use a slightly modified architecture to predict node scores on the line graph representation $\mathcal{G}_{L}$, which directly correspond to the edge scores in $\mathcal{G}_{H}$ (see Figure 3). Following is the formulation of edge and node updates in a generic message passing formulation, as well as the final scoring function based on node representation on $\mathcal{G}_{L}$. In a first step, updated edge representations are calculated using a concatenated vector containing the edge attributes, describing the scattering behavior of the triplet-candidate, together with the node attributes $\bm{v}_{i}^{(0)}$ and $\bm{v}_{j}^{(0)}$ of adjacent nodes according to
@@
Here $\phi_{R,1}$ is a multilayer feed-forward (FF) network mapping the concatenated input to a fixed size representation with $\phi_{R,1}:\mathbb{R}^{2d_{\text{node}}+d_{\text{edge}}}\rightarrow\mathbb{R}^{d_{\text{hidden}}}$.
In a subsequent step, all node features of $\mathcal{G}_{L}$ are updated in an aggregation step using the current node representation and edge information aggregated from the direct neighborhood of the respective node $v_{i}$ according to
$\bm{v}_{i}^{\;(1)}=\phi_{O}\left(\bm{v}_{i}^{\;(0)},\max_{j\in\mathcal{N}_{i}}\left(\bm{e}_{ij}^{\;(1)}\right)\right),$ (10)
and transformed using a learned mapping, parameterized by a feed-forward network, with $\phi_{O}:\mathbb{R}^{d_{hidden}}\rightarrow\mathbb{R}^{d_{\text{hidden}}}$. Given the dynamic number of incoming edges for each graph vertex, either due to the changing numbers of hit readouts in subsequent layers of $\mathcal{G}_{H}$ or the required generalization ability to different particle densities of the proton beam, we use a node-agnostic max-aggregation scheme, providing empirically better generalization abilities compared to non-node-agnostic aggregation schemes (Joshi et al., 2021).
To predict for each edge in $\mathcal{G}_{H}$ an edge score, we perform a final reasoning step, transforming the node embeddings of each node in $\mathcal{G}_{L}$ into a single scalar output using a feed-forward network $\phi_{R,2}$:
@@
We generate the final edge scores $s_{ij}\in[0,1]$ using a sigmoid activation function as proposed in Heintz et al. (2020); DeZoort et al. (2021). We augment the multi-layer feed-forward networks using batch normalization (Ioffe & Szegedy, 2015), to mitigate the risk of vanishing gradients and smoothen the optimization landscape (Santurkar et al., 2018), providing better training and inference performance.
@@ 4.3 Track Construction as a Differentiable Assignment Problem
Generating discrete track candidates from the continuous edge-score predictions requires the application of a discrete (non-differentiable) algorithm that matches the tracks with the lowest total cost. Optimally, this would require solving a multidimensional assignment problem (MAP) to minimize the total cost of all assignments over the entire detector geometry. This exact definition is, however, impractical in real-world usage since solving a MAP is NP-hard (Pierskalla, 1968), requiring approximate solutions. In the related literature, different optimization procedures, such as DBSCAN, union-find or partitioning algorithms, are used frequently (DeZoort et al., 2021; Ju et al., 2021; Lieret & DeZoort, 2023) independently of the training procedure, to construct disconnected tracks during inference. In this work, we rely on a relaxation of the original MAP formulation to a layer-wise linear sum assignment problem (LSAP), finding a solution $y(\bm{\mathcal{C}})$ in form of discrete assignments that minimizes
$\begin{split}y(\bm{\mathcal{C}}):\underset{y\in\mathcal{Y}}{\quad{\arg}\min}&\quad\sum_{(i,j)\in\mathcal{E}}y_{ij}c_{ij}\\ \text{s.t.}&\quad\sum_{i\in\mathcal{V}_{S}}y_{ij}=1,\quad j\in\mathcal{V}_{T}\\ &\quad\sum_{j\in\mathcal{V}_{T}}y_{ij}\leq 1,\quad i\in\mathcal{V}_{S}\\ &\quad y_{ij}\in\{0,1\},\quad i,j\in\mathcal{V}_{T}\times\mathcal{V}_{S}\end{split}$ (12)
for every detector layer. Here, $c_{ij}\in\bm{\mathcal{C}}$ is the assignment cost for the edge combination $i,j$, where $i\in\mathcal{V}_{S}$ is a node from the set of source nodes, defined by all particle hits in the detector layer $l$, and $j\in\mathcal{V}_{T}$ is a node from the set of target nodes in the adjacent detector layer l1l-1 respectively. Further, $y_{ij}$ is the respective assignment policy defining whether the edge should be assigned under the reconstruction policy. We provide, as described below, additional gradient information w.r.t. $\bm{\mathcal{C}}$ through a linearized representation of the solver mapping (Vlastelica et al., 2020), enabling the efficient optimization of the combined tracking algorithm, directly minimizing the assignment error. While the results in Section 5.1 demonstrate good performance for the choice of the LSA assignment, we want to note here that an increasing amount of detector noise and particle scattering can significantly influence the performance due to the required assignment without any notion of noise.
@@
[…]
@@
[…]
@@
The usage of discrete assignment does not provide any confidence scores, thus task losses of correctly assigned edges are always exactly zero, even for marginal differences between cost elements, potentially having a noticeable impact on the generalization performance (Rolínek et al., 2020b; a). To enforce larger margins between predicted costs, Rolínek et al. (2020b; a) introduced ground-truth-induced margins where a negative or positive penalty is added on the predicted costs based on the ground truth labels. Later, Sahoo et al. (2023) introduced noise-induced margins, removing the previous ground-truth dependence by adding random noise to the predicted costs. However, for the large number of edges in our use case of particle tracking, we found both methods to be highly unstable, even for small margin factors.
@@ 4.4 Network Optimization
$\mathcal{L}(\{y_{n}\}_{1:N},\{f(\mathcal{G}_{n})\}_{1:N})=\frac{1}{N}\sum_{n=1}^{N}\frac{1}{|\mathcal{E}_{n}|}\sum_{(i,j)\in\mathcal{E}_{n}}w_{ij}\cdot\left(y_{ij}\cdot(1-\hat{y}_{ij})+(1-y_{ij})\cdot\hat{y}_{ij}\right),$ (15)
@@ 5 Experimental Results and Analysis
For the studies reported in this work, we rely on MC simulation of detector readout data, which we generate using GATE (Jan et al., 2004; 2011) version 9.2 based on Geant4 (version 11.0.0) (Agostinelli et al., 2003; Allison et al., 2006; 2016). We generate different datasets for training (100,000 particles) and validation (5,000 particles), using a 230 MeV pencil beam as a beam source with water phantoms of various thicknesses between particle beam and detector. For comparability purposes, the test set, containing 10,000 particles in total, is taken from Kortus et al. (2022). A detailed listing of all data sources is provided in Appendix F. We generate for the training and validation simulations hit- and line graphs with 100 primaries per frame ($p^{+}/F$), according to the description in Section 4.1, which we consider the expected target density for the constructed detector. Further, we generate hit graphs for the test set with 50, 100 and 150 $p^{+}/F$ to assess the generalization ability to unseen densities.
@@
As baseline models, providing reference values for quantifying the track reconstruction performance of the proposed architectures, we select both a traditional iterative track follower algorithm (Pettersen et al., 2020) [^7] The source code for the track follower proposed by Pettersen et al. is provided as a software component in the Digital Tracking Calorimeter Toolkit (https://github.com/HelgeEgil/DigitalTrackingCalorimeterToolkit), and a reinforcement learning (RL) based tracking algorithm (Kortus et al., 2023). The first baseline finds suitable tracks by iteratively searching for track candidates minimizing the total scattering angle. Similarly, the RL-based algorithm aims to learn a reconstruction policy functioning as a learned heuristic to the algorithm by (Pettersen et al., 2020) by iteratively maximizing the expected likelihood of observing the scattering behavior determined by MCS during training. Similar to our line-graph parameterization in Section 4.1, Kortus et al. (2023) augment the simpler hit-graph features using segment pair information as an inductive bias for tracking. Yet, due to the sequential nature of RL, the pairwise information is queried only for the current partial track candidates without explicitly building the full line-graph (see Kortus et al. (2023) for more details).
@@ 5.1 Training and Inference Performance
Figure 4: True positive and false positive rate of assignments together with purity and efficiency of track reconstruction, evaluated for predict-and-track and predict-then-track as a function of training steps.
@@
Additional performance results are provided in Table 1. Here, the performance of the trained models is compared for different phantom and particle density configurations. In addition, the performance results of a traditional track follower procedure, which was developed for the particular use-case of particle tracking in the DTC prototype, are provided as a baseline.
Table 1: Reconstruction performance (purity $p$ and efficiency $\epsilon$) for 100-200 mm water phantoms and 50-150 primaries per frame $p^{+}/F$. Results marked with * use a layer-wise reconstruction scheme, with the initial transition generated using ground truth.($\uparrow$: higher is better; $\downarrow$ lower is better)
@@
[…]
Similar to the results provided in Figure 4, all model configurations, trained on the task- and the prediction-loss, demonstrate near identical reconstruction purity and efficiency over all tested phantom and particle configurations. Analogous to Vlastelica et al. (2020), we find that the exact choice of $\lambda$ is not critical for the reconstruction quality and only impacts the trainings- and convergence speed with larger $\lambda$ (ref. Figure 4). Further, the graph neural networks outperform the baseline algorithm in almost all performance metrics over all configurations, with only the configuration of 150 $p^{+}/F$ and 100 mm water phantom as a slight outlier in this tendency. Here, the track follower and RL-based tracker demonstrate a higher reconstruction efficiency, while the reconstruction purity is still higher for the GNN architectures.
@@ 5.2 Evaluation of Local and Global Loss Landscapes
To analyze and characterize the training and inference behavior of the end-to-end and two-step tracking approaches and gain an understanding of the effects of end-to-end optimization and its hyperparameters, as well as an intuition regarding the usability and effectiveness of the approach in an E2E-differentiable processing pipeline (as discussed in Section 3.2), we visualize (Li et al., 2018) and characterize the loss landscape structure, closely following the methodology and taxonomy by Yang et al. (2021). We specifically focus on the global connectivity of the loss landscape, as it allows us to infer conclusions about the optimization of particle tracking in an isolated setting as well as an integrated operation in a processing pipeline, also considering gradient information from downstream tasks. We further augment the evaluation by additional analysis of functional similarities (Klabunde et al., 2023) of the optimized models, allowing us to quantify the risk of predictive instabilities of the reconstruction network as well as potential opportunities for E2E fine-tuning, gained by diverse but connected minima. All used algorithms are detailed in Section 3.3. Additional implementation details are listed in Appendix E.